OpenStudy (anonymous):

How do you find tension? I want to get a better general understanding of how to work with it using newtonian mechanics. In my case I am given that there are 2 bricks, on is on top each other vertically connected by a string. Force of 200N applied vertically and it accelerates at 1.2 ms. How do I understand in how to find the tension of the string? I tried applying various forces but the answer doesn't seem to match.

OpenStudy (shamim):

Unit of acceleration is ms^-2

OpenStudy (shamim):

200-mg=ma 200-m*9.8=m*1.2